OpenStudy (anonymous):

Using the properties of exponents and radicals, design at least three different equivalent forms of x². You must show how each one can be simplified back to x² in two or more steps. Stretch your mind and get creative! Keep in mind that something too simple, like x • x would not be acceptable since it takes only one step to convert it to x².

OpenStudy (anonymous):

i really have no idea what to do someone help me!

OpenStudy (anonymous):

\[\frac{ x ^{2}\times x ^{2} }{ ^{x ^{2}} }\]

OpenStudy (anonymous):

thank you but now how do i do this?

OpenStudy (anonymous):

\[x ^{2}+x ^{2}=x ^{4}\]

OpenStudy (anonymous):

oh yea now i get it thank you now i have to find 2 more problems

OpenStudy (anonymous):

\[x ^{4}/x ^{2}=x ^{2}\]

OpenStudy (anonymous):

you can follow that same principle and just add more x's

OpenStudy (anonymous):

x(x^5)/x(x^3)

OpenStudy (anonymous):

ok so i would get x^5 / x3 right?

OpenStudy (anonymous):

no that would be x^6/x^4 which is x^2

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

when you have to terms with the same base and you multiply them you just add the exponent. when you divide two terms with the same base you can just subtract the exponent

OpenStudy (anonymous):

ok thank you for explaining that to me

OpenStudy (anonymous):

ok so is this right (x^-1 *x^-1)X^1= x^1 *X^1 = x^2